OpenStudy (anonymous):

Juice has 30% oj and another has 10% oj. How many liters of each is needed to get 20L at 16% oj?

OpenStudy (anonymous):

Juice = 0.7W + 0.3R Other = 0.9W + 0.1R Water = W Oj = R (real orange juice) you want X(0.7W + 0.3R) + Y (0.9W + 0.1R) = 20, where X & Y are the litre volumes of each kind of drink mixture. you need a total R = 0.16*20 = 3.2 you need a total W = 0.84*20 = 16.8 using the R requirement: 0.3X + 0.1Y = 3.2 using the W requirement: 0.7X + 0.9Y = 16.8 solve the simultaneous equations to find X & Y